# Multi-task Planar Reconstruction with Feature Warping Guidance

### v1

@@ Abstract
Piece-wise planar 3D reconstruction simultaneously segments plane instances and recovers their 3D plane parameters from an image, which is particularly useful for indoor or man-made environments. Efficient reconstruction of 3D planes coupled with semantic predictions offers advantages for a wide range of applications requiring scene understanding and concurrent spatial mapping. However, most existing planar reconstruction models either neglect semantic predictions or do not run efficiently enough for real-time applications. We introduce SoloPlanes, a real-time planar reconstruction model based on a modified instance segmentation architecture which simultaneously predicts semantics for each plane instance, along with plane parameters and piece-wise plane instance masks. By providing multi-view guidance in feature space, we achieve an improvement in instance mask segmentation despite only warping plane features due to the nature of feature sharing in multi-task learning. Our model simultaneously predicts semantics using single images at inference time, while achieving real-time predictions at 43 FPS. The code will be released post-publication.
@@ 1 INTRODUCTION
Figure 1: Comparison of SoloPlanes output with ground truth (GT). 3D projections using predicted plane parameters (left) and GT depth (right). Textures uses RGB (top), predicted semantics (bottom left), and GT (bottom right).
Several methods have been proposed to use a deep learning approach to recover planes of indoor scenes from a single image (Liu et al., 2018b, ; Yu et al.,, 2019; Liu et al., 2018a, ; Xie et al., 2022b, ; Xie et al.,, 2021). While existing works have made strides in predicting piece-wise instance masks and plane parameters, they often ignore the added information from scene semantics. Incorporating semantics provides an added layer of scene understanding, which can be useful in many applications. For instance, the semantic label for a planar surface can help a service robot to determine the correct behaviour (e.g. mopping floor vs. wiping table), or AR/VR experiences could offer semantics-dependent retexturization. Some models predict semantics along with plane parameters but are often too computationally intensive to meet the real-time requirements of practical applications (Liu et al.,, 2022).
@@
We introduce multi-view guidance in feature space by warping plane features from neighbour to source view, decoding, then transforming the decoded plane parameters to the source view camera view for comparison with ground truth data during training. This additional warping guidance for plane features positively impacts the learning of segmentation masks, particularly when using a more limited dataset, while only requiring a single view at inference time.
To the best of our knowledge, this work is the first empirical demonstration of cross-task improvement using multi-view guidance by feature warping, with particular relevance in the case where ground truth data may be incomplete across neighboring views. This approach may be a helpful method for other multi-task models limited in some forms of ground truth training data. Additionally, we introduce a single-image planar reconstruction model that can concurrently predict semantics for planar segments while achieving the best efficiency of all known planar reconstruction methods at a processing speed of 43 FPS.
@@ 2 RELATED WORK
Planar Reconstruction Early works in planar reconstruction using a single image predicted a set number of planes per scene without using ground truth plane annotations by employing a plane structure-induced loss (Yang and Zhou,, 2018). Another early end-to-end planar reconstruction network from a single image is PlaneNet, which uses separate branches for plane parameter, mask, and non-planar area depth estimation (Liu et al., 2018b, ). Two major subsequent models serve as the foundation for several later works. PlaneRCNN is an extension of the two-stage instance segmentation model, Mask-RCNN (He et al.,, 2018), and predicts the plane instance normal and depth map, then jointly process plane parameters along with segmentation masks through a refinement module (Liu et al., 2018a, ). PlaneAE predicts per-pixel parameters and associative embeddings and employs efficient mean clustering to group the pixel embeddings to plane instances (Yu et al.,, 2019). Additional contributions in this area include post-processing refinement networks that enforce interplane relationships, via predicting the contact line or geometric relations between adjacent planes (Qian and Furukawa,, 2020). More recent works follow the approach of using an instance segmentation model base. PlaneSegNet is based on a real-time instance segmentation architecture and introduces an efficient Non-Maximum Suppression (NMS) technique to reduce redundant proposals (Xie et al.,, 2021). PlaneRecNet predicts per-pixel depth and plane segmentation masks, then use classical methods like PCA or RANSAC to recover plane parameters (Xie et al., 2022b, ). A number of single-image plane reconstruction models use some form of instance segmentation baseline. However, most single-image models focus solely on spatial parameters and largely ignore the task of recovering semantics.
@@
Multi-view approaches The task of predicting 3D plane parameters from a single image is inherently ambiguous and challenging. Thus, several works have incorporated multi-view information, either as a loss guidance or by using multiple image inputs at inference time. PlanarRecon (Xie et al., 2022a, ) is a real-time model using multiple image frames which makes predictions directly in 3D by using a coarse-to-fine approach for building a sparse feature volume, then clustering occupied voxels for instance planes, and uses a tracking and fusion module to get a global plane representation. PlaneMVS (Liu et al.,, 2022) is the first to apply a deep multi-view stereo (MVS) approach to plane parameters, Although it achieves state-of-the-art results and also predicts class semantics, it is less computationally efficient due to the use of 3D convolutions and requires generation of plane hypotheses. PlaneRCNN incorporates a multi-view warping loss module that enforces consistency with nearby views by projecting the predictions to 3D and calculating the distance after transforming to the same camera coordinates (Liu et al., 2018a, ). Unlike our approach, their warping module is applied directly on the predictions rather than in feature space. Another work enhances the PlaneAE model with multi-view regularization by warping the plane embedding feature maps and using associative embeddings from multiple views to improve the final instance segmentation (Xi and Chen,, 2019).
@@
Feature Warping Feature warping is commonly done in deep Multi-View Stereo (MVS) approaches, as it was found that creating the cost volume using features is as effective for artificial neural networks and more computationally efficient due to reduced size (Im et al.,, 2019; Yao et al.,, 2018). While some approaches use a similarity function on the features, others simply concatenate the warped feature with the original and let the model learn the relation rather than calculate an explicit cost volume (Chen et al.,, 2020; Yao et al.,, 2018). The latter approach is used by PlaneMVS to construct a Feature/Cost volume, which is then processed by a 3D CNN to get the plane parameters. Deep MVS methods are more commonly used for depth estimation, and their application to plane parameter estimation is relatively novel. The work from Ştefan Săftescu and Newman, (2020) warp feature maps between views for correcting 3D reconstructions, enabling the avoidance of memory-intensive 3D calculations. Other research suggests that calculating a feature error between frames is more robust than a photometric error (Guo et al.,, 2021). However, this cannot directly be applied to plane reconstruction, as the plane features contain information in different camera views when considering a video dataset. Takanori et al. use multi-frame attention via feature warping for the task of drone crowd tracking (Asanomi et al.,, 2023). Ding et al. take MVS as a feature matching task and use a Transformer model to aggregate long-range global context using warped feature maps (Ding et al.,, 2022).
@@
In order to ensure differentiability, the warp to another view using depth values and camera parameters must be backprojected using bilinear interpolation. Most existing works involving feature warping do not specifically deal with plane features, which require transformation to the correct view when decoded. Work from Ştefan Săftescu and Newman, (2020) addresses the view transformation of depth features. Additionally, the majority of planar reconstruction models do not offer semantic predictions for the scene.
@@
The majority of existing works are primarily focused on the geometric accuracy of planes without holistically addressing the more practical requirements of speed and semantic understanding of planar scenes. Our work aims to fill this gap by offering a unified framework for semantic planar reconstruction. We improve data efficiency during training and achieve cross-task improvement using multi-view guidance for plane features, while maintaining an inference speed that is suitable for real-time applications.
@@ 3 METHOD
Our objective is to develop a real-time framework for the task of semantic planar reconstruction. This section is organized as follows: Section 3.1 provides details of the framework, Section 3.2 elaborates on the loss terms, and Section 3.3 introduces our multi-view guidance.
@@ 3.1 Framework
Base Architecture Our framework is built on a light version of the SOLOv2 instance segmentation model (Wang et al., 2020b, ) using a ResNet-50 backbone (He et al.,, 2015). SOLOv2 is a single-stage instance segmentation model that predicts instance masks and labels based on their spatial location. It achieves an execution speed of around 31 FPS using a single V100 GPU card (Wang et al., 2020b, ). The model employs dynamic convolution to generate the final segmentation mask, leveraging multi-scale features from the Feature Pyramid Network (FPN) (Lin et al.,, 2017). Each level of the FPN output features are used to predict mask kernels and class semantics, with the features reshaped to square grids of varying sizes, with each responsible for predictions at a different scale. Each grid location predicts a kernel and semantic category scores. The mask feature is obtained through feature fusion of the first four levels of the FPN outputs via bilinear upsampling and convolution layers, and the final segmentation masks are obtained via convolution using the predicted kernels, with redundant masks suppressed using matrix Non-Maximum Supression (NMS) (Wang et al., 2020b, ). The mask and kernel features receive spatial awareness information from concatenated normalized coordinate, a method from Liu et al., 2018c .
Figure 2: Simplified overview of SoloPlane architecture.
Modifications We extend the base architecture by introducing a plane feature branch that fuses the first two levels of the feature map, along with a plane prediction head that outputs per-pixel plane parameters via a convolution layer. This prediction is supervised by a set of loss functions that leverage geometrical constraints and ground truth depth information (detailed in Section 3.2). The original architecture predicts the kernels and semantic categories using all five feature map levels of the FPN. Based on the findings of Chen et al., (2021), a divide-and-conquer approach is more crucial than leveraging multi-scale features for task-specific predictions, we experimented with using different feature levels and found that using fewer feature levels not only maintained comparable performance in multi-task planar segmentation but also improved the overall efficiency of the model.
Our final architecture takes a single RGB image, $I\in\mathbb{R}^{H\times W\times C}$, as input during inference, and outputs an arbitrary number of plane instance masks along with instance level semantics and per-pixel plane parameters. We obtain the final result by pooling per-pixel parameter prediction using the predicted masks, and retaining per-pixel predictions in areas without a plane instance. The model is trained on the large-scale public ScanNet dataset containing indoor scenes from Dai et al., (2017), supplemented with ground truth plane annotations from Liu et al., 2018a .
@@ 3.2 Losses
Mask & Category We retain the original loss functions from Wang et al., 2020b for mask and category predictions. The Dice Loss, $L_{M}$, guides mask prediction with the original loss weight $w_{M}=3$, and focal loss, $L_{C}$, for semantic category prediction (Lin et al.,, 2018). For full details, we refer readers to (Wang et al., 2020a, ). In order to address class imbalances due to dominating negative samples, we modified $L_{C}$ to only consider grid locations containing an instance.
@@
Gradient Weighting We added gradient edge weighting as a model variation, weighting $L_{depth}$ and $L_{geom}$ to emphasize learning at edges, areas which are typically more difficult to learn. We choose to use the gradient of the image, $G\in\mathbb{R}^{H\times W}$ rather than depth. Despite more noise at non-edge areas, it can capture more plane edges as some plane instances can have the same depth but still represent different surfaces (e.g. picture frame on a wall). This addition results in cross-task improvements for segmentation mask prediction in the case of the multi-view model (see Section 3.3).
@@ 3.3 Multiview Plane Feature Guidance
In this section, we introduce our multi-view guidance approach, depicted in Fig 3. We take neighbouring image pairs, which we denote by source and neighbouring view $(I_{s},I_{n})$, and extract the corresponding 2D features. The two finest pyramid feature maps are fused to generate plane features $f\in\mathbb{R}^{\frac{1}{4}H\times\frac{1}{4}W\times C}$. We backproject the neighbouring feature $f_{N}$ to the corresponding location of the source view using bilinear interpolation. This process uses the ground truth depth, intrinsic parameters, and the relative transform between the views to obtain the warped 2D coordinates, from which we obtain the outprojection mask. We then decode the warped neighbouring feature $\hat{f}_{N}$ with the plane prediction head to get the corresponding plane parameters. It is important to note that $\hat{f}_{N}$ contains plane information of the neighbouring view, under the camera coordinates of $I_{n}$. Therefore, we transform the decoded plane parameters to the source view’s camera coordinates before comparing to ground truth. This transformation is given by:
@@
where $(R,\mathbf{t})$ represents the rotation matrix and translation vector from neighbour to source view, and $(\mathbf{n_{n}},d_{n})$ are the normal and offset in the neighbouring view. We then calculate an additional plane loss $L_{P}$ using the transformed plane parameters decoded from the warped feature, excluding from the loss areas that are occluded or fall outside of the 2D image coordinates using the outprojection mask.
@@ 3.4 Instance Plane Soft-Pooling
To obtain the final instance level plane parameters, we use a soft-pooling technique which only considers per-pixel parameters within the area of the predicted instance. We found that restricting the pooling to this binary area yielded better results compared to using soft-pooling across all pixel locations. We opted to not use an instance level plane loss as it negatively impacted the learning of mask segmentation. We generate a binary segmentation mask by applying a threshold to the predicted soft mask, denoted as $m^{*}\in[0,1]$. The instance level parameter can be retrieved by

### v2

@@ Abstract
Piece-wise planar 3D reconstruction simultaneously segments plane instances and recovers their 3D plane parameters from an image, which is particularly useful for indoor or man-made environments. Efficient reconstruction of 3D planes coupled with semantic predictions offers advantages for a wide range of applications requiring scene understanding and concurrent spatial mapping. However, most existing planar reconstruction models either neglect semantic predictions or do not run efficiently enough for real-time applications. We introduce SOLOPlanes, a real-time planar reconstruction model based on a modified instance segmentation architecture which simultaneously predicts semantics for each plane instance, along with plane parameters and piece-wise plane instance masks. We achieve an improvement in instance mask segmentation by including multi-view guidance for plane predictions in the training process. This cross-task improvement, training for plane prediction but improving the mask segmentation, is due to the nature of feature sharing in multi-task learning. Our model simultaneously predicts semantics using single images at inference time, while achieving real-time predictions at 43 FPS. Code is available at: https://github.com/fraunhoferhhi/SOLOPlanes.
@@ 1 INTRODUCTION
Figure 1: Comparison of SOLOPlanes output with ground truth (GT). 3D projections using predicted plane parameters (left) and GT depth (right). Textures use RGB (top), predicted semantics (bottom left), and GT semantics (bottom right).
Several methods have been proposed to use deep learning to recover planes of indoor scenes from a single image (Liu et al., 2018a, ; Yu et al.,, 2019; Liu et al.,, 2019; Xie et al., 2021b, ; Xie et al., 2021a, ). While existing works have made strides in predicting piece-wise instance masks and plane parameters, they often ignore the added information from scene semantics. Incorporating semantics provides an added layer of scene understanding, which can be useful in many applications. For instance, the semantic label for a planar surface can help a service robot in determining the correct behaviour (e.g. mopping floor vs. wiping table), or AR/VR experiences could offer semantics-dependent retexturization. Some models predict semantics along with plane parameters but are often too computationally intensive to meet the real-time requirements of practical applications (Liu et al.,, 2022).
@@
Our aim was to create a data-efficient model with improved run-time efficiency compared to existing models for planar reconstruction with semantics. We achieve the desired outcome via our model, SOLOPlanes (SOLOP), where we make use of multi-view guidance for improved data usage when acceptable ground truth plane segments differ across views, and made adjustments to the base architecture for improved efficiency. Multi-view warping is done in feature space, by warping plane features from neighbour to source view, decoding, then transforming the decoded plane parameters to the source view camera view for comparison with ground truth data during training. This additional warping guidance for plane features positively impacts the learning of segmentation masks, particularly when using a more limited dataset, while only requiring a single view at inference time.
@@
Our contributions include the following:
  1. 1.
    An empirical demonstration of cross-task improvement using multi-view guidance by feature warping, with particular relevance in cases where ground truth data may be incomplete across neighboring views.
  2. 2.
    A single-image planar reconstruction model, that can concurrently predict semantics for planar segments while achieving the best efficiency compared to other known planar reconstruction methods at a processing speed of 43 FPS.
Our approach may be a helpful method for other multi-task models limited in some forms of ground truth training data. The efficiency of the model makes it suitable for a range of real-world applications.
@@ 2 RELATED WORK
Planar Reconstruction Early works in planar reconstruction using a single image predicted a set number of planes per scene without using ground truth plane annotations by employing a plane structure-induced loss (Yang and Zhou,, 2018). Another early end-to-end planar reconstruction network from a single image is PlaneNet, which uses separate branches for plane parameter, mask, and non-planar area depth estimation (Liu et al., 2018a, ). Two major subsequent models serve as the foundation for several later works. PlaneRCNN is an extension of the two-stage instance segmentation model, Mask-RCNN (He et al.,, 2017), and predicts the plane instance normal and depth map, then jointly process plane parameters along with segmentation masks through a refinement module (Liu et al.,, 2019). PlaneAE predicts per-pixel parameters and associative embeddings and employs efficient mean clustering to group the pixel embeddings to plane instances (Yu et al.,, 2019). PlaneTR uses geometric guidance by generating and tokenizing line segments, giving the input additional structural information (Tan et al.,, 2021). Additional contributions in this area include post-processing refinement networks that enforce interplane relationships, via predicting the contact line or geometric relations between adjacent planes (Qian and Furukawa,, 2020). More recent works follow the approach of using an instance segmentation model base. PlaneSegNet is based on a real-time instance segmentation architecture and introduces an efficient Non-Maximum Suppression (NMS) technique to reduce redundant proposals (Xie et al., 2021a, ). PlaneRecNet predicts per-pixel depth and plane segmentation masks, then use classical methods like PCA or RANSAC to recover plane parameters (Xie et al., 2021b, ). A number of single-image plane reconstruction models use some form of instance segmentation baseline. However, most single-image models focus solely on spatial parameters and largely ignore the task of recovering semantics.
@@
Multi-view approaches The task of predicting 3D plane parameters from a single image is inherently ambiguous and challenging. Thus, several works have incorporated multi-view information, either as a loss guidance or by using multiple image inputs at inference time. PlanarRecon (Xie et al.,, 2022) is a real-time model using multiple image frames which makes predictions directly in 3D by using a coarse-to-fine approach for building a sparse feature volume, then clustering occupied voxels for instance planes, and uses a tracking and fusion module to get a global plane representation. PlaneMVS (Liu et al.,, 2022) is the first to apply a deep multi-view stereo (MVS) approach to plane parameters. Although it achieves state-of-the-art results and also predicts class semantics, it is less computationally efficient due to the use of 3D convolutions and requires generation of plane hypotheses. PlaneRCNN incorporates a multi-view warping loss module that enforces consistency with nearby views by projecting the predictions to 3D and calculating the distance after transforming to the same camera coordinates (Liu et al.,, 2019). Unlike our approach, their warping module is applied directly on the predictions rather than in feature space. Another work enhances the PlaneAE model with multi-view regularization by warping the plane embedding feature maps and using associative embeddings from multiple views to improve the final instance segmentation (Xi and Chen,, 2019).
@@
Feature Warping Feature warping is commonly done in deep Multi-View Stereo (MVS) approaches, as it was found that creating the cost volume using features is as effective for artificial neural networks and more computationally efficient due to reduced size (Im et al.,, 2019; Yao et al.,, 2018). While some approaches use a similarity function on the features, others simply concatenate the warped feature with the original and let the model learn the relation rather than calculate an explicit cost volume (Chen et al.,, 2020; Yao et al.,, 2018). The latter approach is used by PlaneMVS to construct a Feature/Cost volume, which is then processed by a 3D CNN to get the plane parameters. Deep MVS methods are more commonly used for depth estimation, and their application to plane parameter estimation is relatively novel. Other research suggests that calculating a feature error between frames is more robust than a photometric error (Guo et al.,, 2021). However, this cannot directly be applied to plane reconstruction, as the plane features contain information in different camera views when considering a video dataset. Takanori et al. use multi-frame attention via feature warping for the task of drone crowd tracking (Asanomi et al.,, 2023). Ding et al. take MVS as a feature matching task and use a Transformer model to aggregate long-range global context using warped feature maps (Ding et al.,, 2022).
@@
In order to ensure differentiability, the warp to another view using depth values and camera parameters must be backprojected using bilinear interpolation. Most existing works involving feature warping do not specifically deal with plane features, which require transformation to the correct view when decoded. Additionally, the majority of planar reconstruction models do not offer semantic predictions for the scene.
@@
The majority of existing works primarily focus on the geometric accuracy of planes without holistically addressing the more practical requirements of speed and semantic understanding of planar scenes. Our work aims to fill this gap by offering a unified framework for semantic planar reconstruction. We improve data efficiency during training and achieve cross-task improvement using multi-view guidance for plane features, while maintaining an inference speed that is suitable for real-time applications.
@@ 3 METHOD
Our objective is to develop a real-time framework for the task of 3D semantic planar reconstruction. This section is organized as follows: Section 3.1 provides details of the framework, Section 3.2 elaborates on the loss terms, and Section 3.3 introduces our multi-view guidance.
@@ 3.1 Framework
Our framework is built on a light version of the SOLOv2 instance segmentation model (Wang et al., 2020b, ) using a ResNet-50 backbone (He et al.,, 2016). SOLOv2 is a single-stage instance segmentation model that predicts instance masks and labels based on their spatial location. It achieves an execution speed of around 31 FPS using a single V100 GPU card (Wang et al., 2020b, ). The model employs dynamic convolution to generate the final segmentation mask, leveraging multi-scale features from the Feature Pyramid Network (FPN) (Lin et al., 2017a, ). Each level of the FPN output features are used to predict mask kernels and class semantics, with the features reshaped to square grids of varying sizes, with each responsible for predictions at a different scale. Each grid location predicts a kernel and semantic category scores. The mask feature is obtained through feature fusion of the first four levels of the FPN outputs via bilinear upsampling and convolution layers, and the final segmentation masks are obtained via convolution using the predicted kernels, with redundant masks suppressed using matrix Non-Maximum Supression (NMS) (Wang et al., 2020b, ). The mask and kernel features receive spatial awareness information from concatenated normalized coordinate, a method from Liu et al., 2018b .
Figure 2: Simplified overview of SOLOPlanes architecture.
We extend the base architecture by introducing a plane feature branch that fuses the first two levels of the feature map, along with a plane prediction head that outputs per-pixel plane parameters via a convolution layer (see Fig. 2). This prediction is supervised by a set of loss functions that leverage geometrical constraints and ground truth depth information (detailed in Section 3.2). The original architecture predicts the kernels and semantic categories using all five feature map levels of the FPN. Based on the findings of Chen et al., (2021), a divide-and-conquer approach is more crucial than leveraging multi-scale features for task-specific predictions, we experimented with using different feature levels and found that using fewer feature levels not only maintained comparable performance in multi-task planar segmentation but also improved the overall efficiency of the model.
Our final architecture takes a single RGB image, $I\in\mathbb{R}^{H\times W\times C}$, as input during inference, and outputs an arbitrary number of plane instance masks along with instance level semantics and per-pixel plane parameters. We obtain the final result by pooling per-pixel parameter prediction using the predicted masks, and retaining per-pixel predictions in areas without a plane instance. The model is trained on the large-scale public ScanNet dataset containing indoor scenes from Dai et al., (2017), supplemented with ground truth plane annotations from Liu et al., (2019).
@@ 3.2 Losses
Mask & Category We retain the original loss functions from Wang et al., 2020b for mask and category predictions. The Dice Loss, $L_{M}$, guides mask prediction with the original loss weight $w_{M}=3$, and focal loss, $L_{C}$, for semantic category prediction (Lin et al., 2017b, ). For full details, we refer readers to (Wang et al., 2020a, ). In order to address class imbalances due to dominating negative samples, we modified $L_{C}$ to only consider grid locations containing an instance.
@@
Gradient Weighting We add gradient edge weighting as a model variation, weighting $L_{depth}$ and $L_{geom}$ to emphasize learning at edges, areas which are typically more difficult to learn. We choose to use the gradient of the image, $G\in\mathbb{R}^{H\times W}$ rather than depth, in order to better capture edges. Despite more noise at non-edge areas, it can capture more plane edges as some plane instances can have the same depth but still represent different surfaces (e.g. picture frame on a wall). This addition results in cross-task improvements for segmentation mask prediction in the case of the multi-view model (see Section 3.3).
@@ 3.3 Multiview Plane Feature Guidance
In this section, we introduce our multi-view guidance approach, depicted in Fig 3. We take neighbouring image pairs, which we denote by source and neighbouring view $(I_{s},I_{n})$, and extract the corresponding 2D features. The two finest pyramid feature maps are fused to generate plane features $f\in\mathbb{R}^{\frac{1}{4}H\times\frac{1}{4}W\times C}$. We backproject the neighbouring feature $f_{N}$ to the corresponding location of the source view using bilinear interpolation. This process uses the ground truth depth, intrinsic parameters, and the relative transform between the views to obtain the warped 2D coordinates, from which we obtain the out-projection mask. We then decode the warped neighbouring feature $\hat{f}_{N}$ with the plane prediction head to get the corresponding plane parameters. It is important to note that $\hat{f}_{N}$ contains plane information of the neighbouring view, under the camera coordinates of $I_{n}$. Therefore, we transform the decoded plane parameters to the source view’s camera coordinates before comparing to ground truth. This transformation is given by:
@@
where $(R,\mathbf{t})$ represents the rotation matrix and translation vector from neighbour to source view, and $(\mathbf{n_{n}},d_{n})$ are the normal and offset in the neighbouring view. We then calculate an additional plane loss $L_{P}$ using the transformed plane parameters decoded from the warped feature, excluding from the loss areas that are occluded or fall outside of the 2D image coordinates using the out-projection mask.
@@ 3.4 Instance Plane Soft-Pooling
To obtain the final instance level plane parameters, we use a soft-pooling technique which only considers per-pixel parameters within the area of the predicted instance. We found that restricting the pooling to this binary area yields better results compared to using soft-pooling across all pixel locations. We opted to not use an instance level plane loss as it negatively impacts the learning of mask segmentation. We generate a binary segmentation mask by applying a threshold to the predicted soft mask, denoted as $m^{*}\in[0,1]$. The instance level parameter can be retrieved by